OpenStudy (anonymous):

URGENT PLEASE HELP c) (1-cos0) (1+1/cos0) simplifies to cos^2 0 d) sin0/1+cos0 + 1+cos0/sin0 simplifies to 2/sin0 @mathslover @MARC_ @master50777

OpenStudy (michele_laino):

please note that: \[(1-\cos \theta)\left( 1+\frac{ 1 }{ \cos \theta }\right)=1+\frac{ 1 }{ \cos \theta }-\cos \theta-1=...\] please continue. And: \[\frac{ \sin \theta }{ 1+ \cos \theta }+\frac{ 1+ \cos \theta }{ \sin \theta }=\] \[=\frac{ (\sin \theta)^{2}+(\cos \theta+1) ^{2}}{ \sin \theta (1+\cos \theta) }=\] \[=\frac{ (\sin \theta)^{2}+(\cos \theta)^{2}+2 \cos \theta +1 }{ \sin \theta(1+\cos \theta) }=...\] please continue